OpenStudy (love_to_love_you):

The table shows the position of a cyclist on a busy road. t (seconds) | 0 | 1.1 | 2.3 | 3.5 | 4 | 4.6 --------------------------------------- s(t) | 0 | 2.5 | 6.4 |12.6 |19.1 | 25.5 What is the instantaneous velocity of the bike at t = 4 seconds?

OpenStudy (love_to_love_you):

@IrishBoy123

OpenStudy (irishboy123):

19.1

OpenStudy (love_to_love_you):

That's not one of the options

OpenStudy (harman.singh):

I think 's' is referring to the distance, not speed.

OpenStudy (harman.singh):

In that case, velocity at 4 secs would be the distance divided by the time. v=d/t v=19.1/4 v=4.775 m/s

OpenStudy (jdoe0001):

hmmm looks like 19.1 from here as well just to understand each other it may not be one of the choices.... so... it may be asking for something else and not clear enough what that is that said, bear in mind that, the choices are there as guidelines, not as a menu to pick from the choices are supposed to reflect your result, not the other way around

OpenStudy (love_to_love_you):

I wish I could medal you all lol, I got the answer right, it was 4.85 m/s

OpenStudy (love_to_love_you):

Thank you!

OpenStudy (irishboy123):

nah mate!! instantaneous velocity over a period?

OpenStudy (harman.singh):

"The table shows the position of a cyclist..." Position usually refers to 'distance'. That's why we do v=d/t to find the velocity.